OpenStudy (hlares):

Prompt: "Prove the set N - N15 is infinite." Where N is the set of Natural Numbers and the 15 is subscript. I have figured out how to prove N is infinite, but I am unsure how to approach this one. Any tips or assistance would be appreciated, thank you!

OpenStudy (ikram002p):

\(N_{15}\) ?

OpenStudy (hlares):

Aye, it's N minus that.

OpenStudy (ikram002p):

oh so it's given that n_15 is finite group is it ?

OpenStudy (hlares):

Aye, it is. The prompt and nothing else on the page mentions having to prove that that part was finite, only that the difference was infinite.

OpenStudy (hlares):

Though I still wondering if I need to use the Lemma that "for every K element N, every subset of N_k is finite" in it though. As said, just not sure how to approach this one. Proofs are not my strong suit aha.

OpenStudy (ikram002p):

well i need to know what N_k stands for, as i felt earlier N_15 is finite so that lemma helps indeed! @ganeshie8 any thoughts ?

ganeshie8 (ganeshie8):

basically you want to show that removing finite number of elements from an infinite set doesn't change the cardinality

OpenStudy (ikram002p):

yes @ganeshie8 but what is N_k ? i mean without that lemma we could never know that N_15 is infinite right ? we need more info about its definition

ganeshie8 (ganeshie8):

I think \(N_{15}\) is a finite set whose elements are first \(15\) natural numbers

ganeshie8 (ganeshie8):

\[N_{15}=\{1,2,3,\ldots,15\}\] @Hlares please confirm if that is correct

OpenStudy (ikram002p):

yeah that make sense !

OpenStudy (hlares):

Aye, that is correct.

ganeshie8 (ganeshie8):

prove \(\mathbb{N}-N_{15}\) is infinite

OpenStudy (hlares):

And thank you. For proving that it will not change the cardinality of N, should I show that N_15 is a subset of N at all?

ganeshie8 (ganeshie8):

would you agree that union of two finite sets is finite ?

ganeshie8 (ganeshie8):

if you do, then you may try proof by contradiction

ganeshie8 (ganeshie8):

(proof by contradiction) Suppose \(S=\mathbb{N}-N_{15}\) is finite. Since union of two finite sets is finite, \(S\cup N_{15}\) must be finite. But \(S\cup N_{15}=\mathbb{N}\) is infinite. Contradiction.

ganeshie8 (ganeshie8):

that proof depends on below facts : 1) \(\mathbb{N}\) is infinite 2) union of two finite sets is finite

OpenStudy (hlares):

Ohhh, that is a lot better way to approach it (and easier than the hodgepodge I was thinking of). Aye, I could write out a proof by contradiction like that then to prove it and dig up the appropriate theorems and lemmas in the book. Thank you so much!

ganeshie8 (ganeshie8):

may i see your proof for \(\mathbb{N}\) infinite ?

ganeshie8 (ganeshie8):

if you have it online already..

OpenStudy (hlares):

You can, just give me a moment to type it up. All of my stuff is written out by hand.

ganeshie8 (ganeshie8):

if it is easy please take a screenshot and attach otherwise it is okay... i vaguely remember a proof using induction but not so sure of the exact method...

OpenStudy (hlares):

Proof: Suppose N is finite. Then, since N =/= (empty set), there exists k such that N is equivalent to N_k. Thus, as N is equivalent to N_k, there exists a one-to-one correspondence and thus function, f, from N_k onto N. Then let n= {f(1), f(2), f(3),..,f(n), f(n+1)}. Then n =/= f(i) for any i element N_k. Thus, f is not onto N. Therefore, N is an infinite set.

OpenStudy (hlares):

Whoop, forgot the first line (before the "Suppose..." one): This is a proof by contradiction.

ganeshie8 (ganeshie8):

Looks good !

ganeshie8 (ganeshie8):

thank you :)

OpenStudy (hlares):

Thanks! Took me longer to work on it than it probably should have, but, as said, proofs are not my strong suit, so I'm just happy when I can get them done ahah.

ganeshie8 (ganeshie8):

try watching it between 30th minute and 40th minute it has a similar proof...

OpenStudy (hlares):

I'll watch that right now then. Thank you for finding it!

ganeshie8 (ganeshie8):

his \(J_n\) is same as your \(N_k\)

OpenStudy (ikram002p):

here another thought. if N is finite then for some k, N=N_k, but there exist m \in N such that m=k+1 which is a contradiction.